### v1

# Identifying the magnetic genes in fully- and partially-ordered V2Al ( Cr, Mn, Fe, Co, Ni) Heusler alloys

Abstract
Multicomponent Heusler alloys exhibit various magnetic properties arising from their diverse atomic compositions and crystal structures. Identifying the general physical principles that govern these behaviors is essential for advancing their potential in spintronic applications. In this work, we combine density functional theory with atomistic Monte Carlo simulations to investigate the magnetic ground states, finite-temperature magnetic transitions, and electronic structures of fully-ordered -, -type, and partially-ordered VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys. We introduce the concept of magnetic genes, defined as V--V triangular motifs connected by the nearest-neighbor (NN) exchange interactions . Within this framework, the magnetic ground states and transition temperatures across the VAl family can be consistently understood. The magnetic order is primarily governed by the NN interactions in the triangular genes, while the transition temperatures are additionally influenced by couplings. Furthermore, the magnetic genes are still proven to be effective in our calculations on partially-ordered V2Al alloys from to -type structures. Our results suggest that the concept of magnetic genes provides a unifying principle for understanding magnetic ordering in V-based Heusler alloys and could serve as a powerful guide for exploring magnetism and designing advanced spintronic materials in a broader class of Heusler systems.
I INTRODUCTION
Heusler alloys, long regarded as prototypical spintronic materials owing to their half-metallic nature, are capable of delivering nearly spin polarization at and even well above room temperature [1, 2, 3, 4, 5, 6]. Consequently, the discovery and design of high-performance Heusler compounds remain a central objective of both experimental and theoretical research [7]. On the experimental side, typical full-Heusler alloys such as Co2CrAl and Co2TiSn have been shown to sustain ferromagnetic order up to room temperature, with Curie temperatures of approximately 335 K and 333 K, respectively [8, 9]. Over the past two decades, theoretical efforts, particularly high-throughput first-principles calculations combined with Monte Carlo or spin-model simulations, have identified dozens of full-Heusler systems exhibiting magnetic anisotropy energies exceeding together with Curie temperatures above [10, 11]. These advances have substantially expanded the library of high-performance magnets and promising high- candidates for spintronic applications [12].
In particular, full-Heusler alloys with the general formula ( transition metals) have attracted considerable attention owing to their rich magnetic properties, shape-memory behavior, and sensitivity to chemical disorder [13, 14, 15]. For instance, Co2VAl exhibits ferromagnetic ordering with a Curie temperature close to room temperature, while its decreases under applied external pressure [16]. In contrast, Mn2VGa shows ferrimagnetic ordering with a significantly higher Curie temperature of 784 K [17]. The inverse-Heusler (-type) compound Mn2CoAl also displays robust ferromagnetism, with a of 720 K at ambient pressure [18], whereas Mn2FeSi, despite sharing the same crystal structure, has an anomalously low of only 67 K [19]. Moreover, both theoretical and experimental studies of -type full-Heusler alloys have revealed complex magnetic behavior. For example, Co2FeSi exhibits a remarkably high Curie temperature of about 1100 K [20], while Fe2VAl remains paramagnetic up to room temperature. Such diverse magnetic orders are generally attributed to variations in stoichiometry and chemical disorder [21], raising fundamental questions about the key factors that govern the magnetic properties of Heusler systems.
Recently, chemical disorder has attracted growing attention as an effective means of tuning magnetic order. Disorder can substantially modify magnetic exchange interactions, influence phase stability, and reshape the overall magnetic state of a material [22, 23, 24]. For example, studies on Ru2Mn ( Sn, Sb, Ge, Si) full-Heusler alloys have demonstrated the strong impact of chemical disorder on their magnetic behavior [22], with Co2FeSn exhibiting enhanced magnetism and pronounced size-dependent disorder effects [25]. Similarly, Smith et al. recently employed DC magnetron sputtering to grow epitaxial V2FeAl thin films [26]. Although V2FeAl shows a relatively high Curie temperature experimentally, its crystal structure appears to be a mixture of the - and -type phases, complicated further by intrinsic disorder [26]. Other works have demonstrated that increasing the degree of chemical disorder leads to a pronounced suppression of spontaneous magnetization and Curie temperature [26, 27]. From a broader physical perspective, the deliberate introduction of chemical disorder not only provides an alternative strategy for tuning magnetism but also establishes a promising platform to investigate the fundamental mechanisms underlying complex magnetic orders in Heusler alloys [24].
Although the magnetic properties of full-Heusler alloys have been extensively investigated experimentally, the underlying mechanisms remain elusive due to their complex atomic compositions, multi-sublattice character, and the presence of chemical disorder. Using first-principles approaches, previous works have systematically evaluated the magnetic exchange couplings for several Heusler alloy families employing frozen-magnon and linear-response techniques [28, 29]. For vanadium-based antiferromagnets, existing studies have primarily concentrated on the ordered -type structure of VAl ( Nb, Ta) [30]. Follow-up investigations extended this work by comparing the phase with the inverse phase across a broader range of compounds. In particular, Kuroda et al. reported that -type V2NbAl and V2TaAl favor antiferromagnetic coupling, whereas in VSi ( Ti, Zr, Hf), the structure often competes with or destabilizes the phase [30]. In addition, large-scale first-principles screenings of the inverse-Heusler chemical space, where Sc, Ti, V, Cr, Mn; Al, Si, As; and spans from Ti to Zn, have identified the necessary conditions for the emergence of semimetallic states and established three Slater–Pauling rules associated with – hybridization [31]. Despite these advances, the systematic experimental synthesis and theoretical calculations of V-based VAl compounds with the complete set of Cr, Mn, Fe, Co, and Ni are still lacking. A comprehensive microscopic understanding of the mechanisms that govern magnetism in both fully- and partially-ordered V-based full-Heusler alloys is therefore essential.
In this work, we investigate the ground-state magnetic orders and electronic structures of VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys using density functional theory. For the fully-ordered -type VAl, V2CrAl and V2MnAl exhibit ferrimagnetic order, whereas the remaining compounds adopt ferromagnetic order. The situation becomes more intricate in the inverse -type VAl. While V2CrAl displays a nearly vanishing net magnetic moment with ferromagnetic order between the two different Wyckoff sites of V atoms, the other four alloys exhibit ferrimagnetic order with antiparallel spin alignment between the two V sublattices. Despite this diversity, the magnetic orders across VAl compounds can be consistently interpreted in terms of the V--V triangular pathway, governed by the NN magnetic exchange coupling . Furthermore, finite-temperature magnetic transitions were systematically investigated through atomistic Monte Carlo simulations, which reveal that the sequence of Curie temperatures among different compositions and lattice types is primarily determined by , with additional contributions from exchange interactions. Importantly, we find that the V--V triangular pathway framework remains valid even in partially-ordered systems from to -type structures. Building on this insight, we propose the concept of “magnetic genes”, defined as the V--V triangular motif connected by NN exchange interactions . We recognize that the magnetic genes could provide a promising concept for exploring the magnetic properties in more Heusler alloys.
II METHODS
To perform the structural relaxation calculations, we use the VASP code [32, 33], which implements the projector-augmented wave (PAW) method [34] and is based on the GGA-PBE generalized gradient approximation [35]. A dense -mesh of 10×10×10 is employed for the structural relaxation calcuations. The plane-wave cutoff was chosen to be 425 eV. The convergence criteria for the total energy and ionic forces were set to and , respectively.
Additionally, we performed spin-polarized density-functional [36, 37] calculations using the Atomic-orbital Based Ab-initio Computation at USTC (ABACUS) package [38, 39], an open-source first-principles platform with advanced ground- and excited-state modules, and interfaces to Wannier90. We employed the SG15 optimized norm-conserving Vanderbilt (ONCV) pseudopotentials [40] and the same exchange correlation [35]. The Kohn-Sham wave functions were expanded using a linear combination of atomic orbitals corresponding to the SG15 ONCV (Version 2.0 standard orbitals) [39]. After convergence tests, the grid cutoff for numerical integrations was set to 100 Ry.
To explore the magnetic properties in V2Al, we employed the TB2J package [41], which is based on the use of Green’s functions. The real-space Heisenberg exchange parameters
where is the on-site exchange splitting and the spin-resolved Green function. The complete set defines a classical Heisenberg Hamiltonian
which was solved with the Vampire atomistic spin-dynamics code [42].
The simulated system of V2Al is a cubic box of edge length , generated by replicating the relaxed face-centered unit cell and containing about spins, which was evolved with a stochastic spin integrator. Temperatures from to in increments were sampled using equilibration steps. The Curie temperature was determined from the inflection point of the magnetization .
| Compound | Source | ||||
| -V2CrAl | 0.901 | -0.742 | 0.026 | 1.085 | This work |
| -V2MnAl | 1.802 | -2.018 | 0.068 | 1.653 | This work |
| -V2MnAl | 1.667 | -1.653 | -0.040 | 1.641 | Jerzy et al. [43] |
| -V2FeAl | 0.537 | 1.917 | -0.021 | 2.970 | This work |
| -V2FeAl | 0.362 | 1.889 | -0.014 | 2.599 | Jerzy et al. [43] |
| -V2CoAl | 1.238 | 1.136 | 0.024 | 3.636 | This work |
| -V2NiAl | 1.249 | 0.265 | -0.005 | 2.757 | This work |
III RESULTS AND DISCUSSION
III.1 Magnetic orders of fully-ordered - and -V2Al
The magnetic ordering in Heusler alloys is strongly influenced by both their atomic composition and crystallographic structure. In this work, we focus on two prototypical lattice types, namely the and structures. As established in previous studies, these two structural variants provide a well-defined platform for examining the role of chemical disorder without altering the overall structural topology [7, 14, 28]. As illustrated in Fig. 1(a), the -type full-Heusler structure (space group , No.225) with the generic formula VAl consists of V atoms occupying the 8 Wyckoff positions, while Al and atoms are located at the 4 and 4 sites, respectively [7]. By contrast, in the -type inverse-Heusler structure (space group , No.216), shown in Fig. 1(b), the atoms occupy the 4 positions that arise from the symmetry breaking of the 8 sites in the -type structure, while the V atoms are shifted to the 4 sites. From a broader perspective, the redistribution of atomic occupations between the 4 and 4 sites, driven by the evolution from to -type structures, highlights the importance of chemical disorder in determining the magnetic properties of these compounds. A detailed analysis of disorder effects will be presented in Sec. III.4.
| Compound | Source | |||||
| -V2CrAl | -0.497 | -1.397 | 1.774 | -0.026 | 0.145 | This work |
| -V2MnAl | -0.820 | 1.466 | 1.610 | 0.007 | 2.264 | This work |
| -V2MnAl | -0.689 | 1.440 | 1.340 | -0.082 | 2.009 | Jerzy et al. [43] |
| -V2FeAl | -0.447 | 2.043 | 1.179 | 0.002 | 2.777 | This work |
| -V2FeAl | -0.790 | 2.180 | 1.699 | -0.120 | 2.950 | Jerzy et al. [43] |
| -V2FeAl | -0.310 | 2.110 | 1.20 | 0.090 | 3.090 | Skaftouros et al. [31] |
| -V2CoAl | -0.404 | 1.801 | 0.556 | 0.043 | 1.996 | This work |
| -V2CoAl | -0.54 | 2.16 | 0.50 | -0.12 | 2.0 | Zhang et al. [44] |
| -V2NiAl | -0.526 | 1.527 | 0.107 | 0.039 | 1.148 | This work |
| -V2NiAl | -0.66 | 1.74 | 0.06 | -0.06 | 1.08 | Zhang et al. [44] |
We begin by examining the effect of atomic composition on the magnetic order in -type VAl ( Cr, Mn, Fe, Co, Ni), where the crystal structure is relatively simple, containing only three inequivalent Wyckoff sites compared to the more complex -type structure. As summarized in Table 1, compounds with Fe, Co, and Ni exhibit ferromagnetic ordering, whereas those with Cr and Mn display ferrimagnetic ordering characterized by antiparallel spin alignment between the and V sublattices. In particular, our calculations show that V2MnAl adopts a ferrimagnetic ground state, with magnetic moments of per V atom and per Mn atom. By contrast, V2FeAl exhibits ferromagnetic ordering with a total magnetic moment of per formula unit, in agreement with previous reports [43]. Notably, the total magnetic moments of the Fe to Ni compounds do not vary monotonically across the series. While the magnetic moments on the V sublattice increase progressively from Fe to Ni, those on the site decrease rapidly, leading to non-monotonic behavior in the total moment. These complex magnetic orders and the non-monotonic variation in total magnetization are primarily attributed to differences in the underlying electronic structures arising from the atomic composition. A detailed discussion of the interplay between electronic structure and magnetic order will be presented in Sec. III.5.
To investigate the influence of lattice type on magnetism, we next examine the -type inverse-Heusler VAl compounds, which possess a more complex crystal structure with four inequivalent Wyckoff sites compared to the -type structure. Interestingly, among these systems, only V2CrAl exhibits ferromagnetic coupling between the two inequivalent V sites, while the magnetic moments at the V-4 and -4 sites align antiferromagnetically, as summarized in Tables 2. This contrasts sharply with the comparatively simple magnetic ordering observed in the -type compounds. More specifically, apart from V2CrAl, all other -type VAl compounds ( = Mn, Fe, Co, Ni) display antiparallel alignment of the spin moments between the V atoms at the 4 and 4 sites. Our calculated magnetic moments for -type VAl ( = Mn, Fe, Co) are in good agreement with previous reports [31, 43, 44]. Furthermore, the half-metallic compounds V2MnAl and V2FeAl exhibit total magnetic moments of 2.26 and 2.78 /f.u., respectively, which are close to the ideal values of 2 and 3 /f.u. predicted by the generalized Slater–Pauling relation of , where is the total valence-electron count [31, 45]. Notably, as varies from Cr to Ni, the local magnetic moment of the site decreases monotonically (from 1.77 to 0.11 ), whereas the total moment evolves in a nonmonotonic manner. These results highlight the intricate and composition-dependent magnetic behaviors of inverse-Heusler VAl compounds, motivating the need for a more general understanding of their underlying mechanisms.
III.2 Microscopic magnetic mechanism in - and -V2Al
To gain deeper insight into the origin of magnetism in these systems, we investigate the magnetic exchange couplings () across different atomic compositions and lattice types. As discussed above, in the relatively simple -type VAl compounds, only those with Cr and Mn exhibit ferrimagnetic order, characterized by anti-parallel spin alignment between the and V atoms. Correspondingly, as shown in Figs. 2(a) and 2(b), the NN exchange coupling () between Mn (Cr) atoms at the 4 site and V atoms at the 8 site is antiferromagnetic type, which has the largest value among all the other magnetic exchange couplings in these materials. In contrast, for Fe, Co, and Ni, the NN coupling between the and V atoms is ferromagnetic, as illustrated in Fig. 2(c). Meanwhile, the exchange interactions between - atoms are significantly weaker and do not play a decisive role in stabilizing the magnetic order. The V-V interactions, however, exhibit more intricate behavior. In V2MnAl, both the NN and next-nearest-neighbor (NNN) V-V couplings, indicated by the gray line in Fig. 2(a), are ferromagnetic. However, the NN V-V coupling in V2CrAl, colored by the gray line, acquires a finite antiferromagnetic contribution, as shown in Fig. 2(b). Taken together, these results indicate that antiferromagnetic V-Mn(Cr)-V exchange pathways form triangular motifs that ultimately stabilize ferromagnetic ordering among the V sublattice. Overall, our exchange-coupling analysis provides a microscopic explanation for the emergence of complex magnetic orders in -type VAl compounds.
Here, we investigate the physical mechanisms underlying the magnetic ordering in the more complex -type VAl compounds. In these type compounds, the magnetic moments of V1 (4 site) and (4 site) are aligned antiparallel, indicating dominant antiferromagnetic exchange coupling. As shown in Figs. 2(d) and 2(e), the NN exchange couplings (highlighted in blue) are strongly antiferromagnetic ( Cr, Mn; couplings for Fe, Co, Ni are omitted for clarity) and represent the primary factor stabilizing ferrimagnetic ordering between the and V1 sublattices. In contrast, the NN exchange coupling between and V2 (4 site), as shown in purple lines in Figs. 2(e) and 2(f), evolves rapidly from antiferromagnetic to ferromagnetic as changes from Cr to Ni. This change of is consistent with the corresponding magnetic ordering from ferrimagnetic for Cr to ferromagnetic for Mn, Fe, Co, and Ni. Notably, although the NN V2-Mn interaction is weakly antiferromagnetic type, the NNN V2-Mn coupling is strongly ferromagnetic, suggesting that V2MnAl lies near a compositional boundary where the magnetic ordering transitions from ferrimagnetic to ferromagnetic with increasing atomic number.
From our analysis of magnetic orders and exchange couplings, a comprehensive understanding of the overall magnetic behavior in VAl across different atomic compositions and lattice types can be established. In -type VAl, the diverse magnetic orders can be rationalized in terms of triangular exchange pathways formed by V--V atoms, as illustrated in the inset of Fig. 2(a). In -type VAl, the underlying mechanism is similar, but an additional type of magnetic exchange coupling is present. Our results indicate that the V1--V2 triangular pathway remains a key determinant of magnetic order in -type compounds, as shown in the inset of Fig. 2(d). Changes in the type of the coupling substantially influence the magnetic ordering in these materials. Moreover, although the direct NN exchange coupling between - atoms ( Mn, Fe, Co, Ni) is antiferromagnetic type in -VAl, the triangular -V1- pathway ultimately stabilizes ferromagnetic ordering due to the intervening antiferromagnetic X-V1 couplings, as illustrated in the inset of Fig. 2(d). In summary, despite variations in atomic composition and lattice type, the site plays a central role in controlling the magnetic properties of V-based VAl compounds through triangular pathways formed by NN V and atoms. We therefore identify these motif as a candidate “magnetic genes”.
III.3 Magnetic transition temperature of - and -V2Al
The microscopic mechanisms of magnetic exchange couplings directly determine the macroscopic magnetic properties. Based on our calculated exchange couplings in VAl, we further simulate the magnetic transition temperatures across this series of Heusler compounds. As shown in Fig. 3(a), in -type VAl, V2MnAl exhibits the highest Curie temperature, reaching approximately 900K, whereas the Curie temperatures of the remaining compounds ( Cr, Fe, Co, Ni) are significantly lower, ranging from 100 K to 300 K. The exceptionally high Curie temperature of V2MnAl originates from the effective ferromagnetic coupling between V atoms mediated by the triangular V-Mn-V pathways, which involve strong antiferromagnetic NN couplings ( meV). This is further reinforced by the direct ferromagnetic V-V interactions. Remarkably, our Monte Carlo simulations reveal that the magnitude of the NN V– exchange coupling predominantly determines the magnetic transition temperature in -type V2Al. Specifically, as shown in Fig. 3(a), decreases from 900 K for V2MnAl to 100 K for V2FeAl, consistent with the corresponding NN exchange couplings , which vary from to 2.0 meV, as illustrated in Fig. 2.
The temperature dependence of magnetic phase transitions in -type VAl compounds differs markedly from that in -type structures. As shown in Fig.3(b), the Curie temperatures in -type compounds vary within a narrower range, from approximately 700 K to 300 K, compared to the broader variation in -type materials. This narrower range reflects the more complex network of magnetic exchange couplings in the inverse-Heusler structure. Specifically, V2CrAl exhibits the highest at about 700 K, while the Curie temperature gradually decreases as changes from Fe to Ni. The magnetic transition temperature of V2MnAl shows a slight anomaly, lying between the values for Co and Ni. Despite the apparent complexity of these trends, they can be rationalized using the proposed microscopic mechanism based on the triangular exchange pathways. As discussed above, the NN exchange coupling involving the -site dominates the magnetic transition behavior. For instance, the relatively large contributes to the higher in V2CrAl. Although is also sizable, the of V2MnAl is significantly lower due to the additional influence of the coupling. In V2CrAl, the ferromagnetic – interactions enhance , whereas in V2MnAl, the coupling is antiferromagnetic, leading to a notable reduction of . In the series Fe, Co, Ni, the gradual decrease of is primarily driven by the weakening of the NN coupling, which diminishes from to 0 meV.
In summary, the magnetic ground states and magnetic transition temperatures of VAl compounds can be rationalized within the framework of the V--V triangular exchange pathway and its associated NN magnetic exchange couplings, applicable to both - and -type structures. For simplicity, the NN exchange coupling primarily governs the magnetic ordering, whereas the combination of and predominantly determines the magnetic transition temperature . Although depends on both the exchange couplings and the local magnetic moments of the constituent atoms, no straightforward correlation between the magnitude of individual magnetic moments and is evident. Nevertheless, the observed trends in the dominant exchange couplings provide a practical framework to anticipate and potentially tune the Curie temperatures in these Heusler alloys.
| Structural ordering | |||||||
| () | () | () | () | () | () | f.u. | |
| 100% | -2.018 | ––– | ––– | 1.802 | 1.802 | 0.068 | 1.653 |
| 75%/25% | -1.863 | 2.936 | -0.872 | 1.184 | 1.432 | 0.033 | 1.472 |
| 50%/50% | -1.976 | 2.731 | -0.926 | 0.630 | 1.474 | 0.016 | 1.719 |
| 25%/75% | -1.593 | 1.582 | -0.595 | 0.111 | 1.338 | 0.028 | 1.679 |
| 100% | ––– | 1.610 | -0.820 | ––– | 1.466 | 0.007 | 2.264 |
III.4 Effect of chemical disorder on magnetism of partially-ordered V2Al
In VAl compounds, the NN magnetic exchange coupling plays a central role in determining both the magnetic ground states and the magnetic transition temperatures. This raises the question of whether this framework remains valid in the presence of chemical disorder. In this section, we investigate the evolution of magnetic order and magnetic transition temperatures under partial chemical disorder, introduced by progressively transforming the -type structure toward the -type structure, as illustrated in Fig. 1(c) and 1(d). The introduction of disorder significantly reduces crystal symmetry, complicating the analysis of magnetic interactions. For clarity, we focus on the -type V2MnAl, which exhibits the highest Curie temperature among its sister compounds. Notably, our calculations show that the magnetic ground states remain robust across different degrees of chemical disorder. As summarized in Tables 3, starting from the fully-ordered -type structure, sequential swapping of Mn atoms in the 4 sites with V atoms in the 8 sites results in only minor reductions in the magnetic moment of V2 (nominal 4 site), which then stabilizes as the structure approaches the -type configuration. Furthermore, despite atom swapping, the spin orientations at the corresponding Wyckoff sites are preserved. The magnetic moments of V atoms on both 4 and 4 sites decrease gradually from the to the -type structure, as do the magnetic moments of Mn atoms on their respective sites. Overall, these results indicate that the intricate network of magnetic exchange couplings in VAl is sufficiently robust to stabilize magnetic order even in the presence of substantial chemical disorder.
The magnetic transition temperature in partially-ordered V2MnAl gradually decreases as the structure evolves from to -type structure. As shown in Fig. 4(a), the Curie temperature decreases from 900 K in the fully-ordered structure to 420 K in the configuration. Within our framework, the NN magnetic exchange couplings and are the primary factors governing the magnetic transition temperature. In the partially-ordered structures studied here, swapping V and atoms directly affects the NN coupling , while other magnetic exchange couplings are also modified. A comparison between the calculated Curie temperatures and the corresponding values in Fig. 4 shows a clear correlation. The magnitude of on the swapped sites closely tracks the variation of . For instance, in 75% -V2MnAl, is reduced to meV, and further decreases to meV ( meV and meV for the other two relevant sites) in 25% -V2MnAl, compared to meV in the fully-ordered structure. The rapid reduction of from to 75% is reflected in the corresponding decrease of from 990 K to 650 K.
Interestingly, within the V2Al series, the trend of magnetic transition temperatures in V2FeAl is opposite to that observed in V2MnAl. As shown in Fig. 5(a), the Curie temperature of V2FeAl gradually decreases from 600 K to 120 K as the structure evolves from - to type. This observation raises the question of which factors primarily govern this anomalous sequence of magnetic transition temperatures. As discussed above, the NN magnetic exchange coupling between V1 and () plays a dominant role in determining . Our calculations indicate that, although the sequence of with increasing chemical order in V2FeAl is reversed relative to V2MnAl, the NN coupling remains a key factor. As shown in Fig. 5(b), the absolute value of decreases systematically from the - to the -type structure, consistent with the observed reduction in . These results confirm the validity of as the primary determinant of the Curie temperature in partially-ordered V2FeAl.
Our calculations on partially chemically disordered VAl indicate that the NN magnetic exchange coupling remains the dominant factor governing the magnetic transition temperature. Notably, during the transition from the to the -type structure, the 4 site is consistently occupied by V2 atoms, ensuring that the V-X-V magnetic triangular pathway is preserved. Consequently, the microscopic magnetic mechanism in chemically disordered VAl is still primarily governed by the V-X-V triangular pathway and the associated NN magnetic exchange couplings .
III.5 Electronic structures of fully-ordered V2Al
Based on the calculated magnetic ground states of -type VAl in Sec. III.1, the observed complex magnetic orders and non-monotonic variation of total magnetic moments with atomic composition are proposed related to their electronic structures. In -type VAl, the relatively small total magnetic moments of V2CrAl and V2MnAl originate from the anti-parallel alignment of spin moments between the V and sublattices. This behavior is clearly reflected in the calculated density of states (DOS). As shown in Fig. 6(a), the total DOS of -type V2Al exhibits a high value at the Fermi level, indicating itinerant magnetism in these compounds [46, 47, 48]. Notably, several compounds display nearly half-metallic character. As shown in Fig. 6(b), Cr and Mn atoms exhibit pronounced occupied states with sharp peaks at approximately -1 eV and -2 eV in the spin-down channel, while their spin-up channel states are minimally occupied. In contrast, for Fe, Co, and Ni, the PDOS of atoms shows significant occupation in the spin-up channel and comparatively low values in the spin-down channel. These electronic features are consistent with the observed transition in magnetic order from ferrimagnetism ( Cr, Mn) to ferromagnetism ( Fe, Co, Ni).
Furthermore, from Fe to Ni, the PDOS of atoms shifts to lower energies in both spin-up and spin-down channels, as shown in Fig. 6(b). This trend is attributed to the increasing number of valence electrons with rising atomic number. The pronounced difference between the spin-up and spin-down PDOS of Fe atoms indicates a large magnetic moment on Fe in V2FeAl, whereas the PDOS of Ni atoms in V2NiAl is nearly symmetric between spin channels, reflecting a significantly reduced magnetic moment. Examination of the V-atom PDOS from Fe to Ni reveals an increase in spin-up states below the Fermi level for V atoms in Co- and Ni-containing compounds, particularly in the energy range from -1 to 0 eV, as shown in Fig. 6(c), while the spin-down PDOS decreases in the same energy range. This behavior correlates with the observed increase in magnetic moments of V atoms from Fe to Ni. In contrast, the PDOS of Al atoms remains largely unchanged across the series in Fig. 6(d), consistent with the negligible magnetic moment on Al atoms. Taken together, the change in magnetic order from Mn to Fe, combined with the opposing trends in magnetic moments of V and atoms from Fe to Ni, accounts for the non-monotonic variation in the total magnetic moment across the V2Al series.
IV CONCLUSION
In this work, we employed density functional theory calculations combined with Monte Carlo simulations based on Heisenberg magnetic exchange couplings to investigate V2Al alloys and introduced the concept of “magnetic genes”. These magnetic genes govern both the magnetic ground states and the magnetic transition temperatures in fully- and partially-ordered compounds. The underlying physical mechanism arises from the modulation of NN magnetic exchange couplings within the V--V triangular pathways, which is influenced by atomic composition, crystal structure, and the degree of chemical disorder. Using the framework of magnetic genes, the magnetic orders can be rationalized in terms of the V--V triangular pathways formed by NN couplings. Furthermore, the magnetic transition temperatures are predominantly determined by both NN and additional magnetic exchange interactions, irrespective of whether the alloy is fully or partially ordered. We anticipate that the identification of magnetic genes in other Heusler alloys may provide a systematic and predictive approach to understanding and tuning their diverse magnetic properties.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants No. 52371174 and No. 12204033), the Fundamental Research Funds for the Central Universities (Grant No. FRF-TP-22-097A1), the State Key Lab of Advanced Metals and Materials (Grant No. 2022Z-13), the Young Elite Scientist Sponsorship Program by BAST (Grant No. BYESS2023301) and the High-performance Computing Platform of University of Science and Technology Beijing.References
- Gupta et al. [2023] S. Gupta, S. Chakraborty, V. Bhasin, S. Pakhira, S. Dan, C. Barreteau, J.-C. Crivello, S. N. Jha, M. Avdeev, J.-M. Greneche, D. Bhattacharyya, E. Alleno, and C. Mazumdar, High spin polarization in the disordered quaternary heusler alloy femnvga, Phys. Rev. B 108, 045137 (2023).
- Venkateswara et al. [2023] Y. Venkateswara, J. Nag, S. S. Samatham, A. K. Patel, P. D. Babu, M. R. Varma, J. Nayak, K. G. Suresh, and A. Alam, FeRhCrSi: Spin semimetal with spin-valve behavior at room temperature, Phys. Rev. B 107, L100401 (2023).
- Nishihaya et al. [2025] S. Nishihaya, M. J. A. Jardine, H. S. Inbar, A. Goswami, J. T. Dong, A. N. Engel, Y.-H. Chang, C. P. Dempsey, M. Hashimoto, D. Lu, N. Marom, and C. J. Palmstrøm, Near-half-metallic state in the half-Heusler PtMnSb film on a III-V substrate, Phys. Rev. Mater. 9, 044403 (2025).
- de Groot et al. [1983] R. A. de Groot, F. M. Mueller, P. G. v. Engen, and K. H. J. Buschow, New class of materials: Half-metallic ferromagnets, Phys. Rev. Lett. 50, 2024 (1983).
- Žutić et al. [2004] I. Žutić, J. Fabian, and S. Das Sarma, Spintronics: Fundamentals and applications, Rev. Mod. Phys. 76, 323 (2004).
- Katsnelson et al. [2008] M. I. Katsnelson, V. Y. Irkhin, L. Chioncel, A. I. Lichtenstein, and R. A. de Groot, Half-metallic ferromagnets: From band structure to many-body effects, Rev. Mod. Phys. 80, 315 (2008).
- Graf et al. [2011] T. Graf, C. Felser, and S. S. Parkin, Simple rules for the understanding of Heusler compounds, Progress in Solid State Chemistry 39, 1 (2011).
- Husmann and Singh [2006] A. Husmann and L. J. Singh, Temperature dependence of the anomalous hall conductivity in the Heusler alloy , Phys. Rev. B 73, 172417 (2006).
- Shigeta et al. [2018] I. Shigeta, Y. Fujimoto, R. Ooka, Y. Nishisako, M. Tsujikawa, R. Y. Umetsu, A. Nomura, K. Yubuta, Y. Miura, T. Kanomata, M. Shirai, J. Gouchi, Y. Uwatoko, and M. Hiroi, Pressure effect on the magnetic properties of the half-metallic Heusler alloy , Phys. Rev. B 97, 104414 (2018).
- Kübler et al. [2007] J. Kübler, G. H. Fecher, and C. Felser, Understanding the trend in the curie temperatures of -based Heusler compounds: Ab initio calculations, Phys. Rev. B 76, 024414 (2007).
- Fortunato et al. [2024] N. M. Fortunato, X. Li, S. Schönecker, R. Xie, A. Taubel, F. Scheibel, I. Opahle, O. Gutfleisch, and H. Zhang, High-throughput screening of all--metal Heusler alloys for magnetocaloric applications, Chemistry of Materials 36, 6765 (2024).
- Marathe and Herper [2023] M. Marathe and H. C. Herper, Exploration of all- Heusler alloys for permanent magnets: An ab initio based high-throughput study, Phys. Rev. B 107, 174402 (2023).
- Guo et al. [2025] Y. Guo, C. Guo, Y. Xu, B. Zheng, X. Ni, and F. Tian, Machine learning investigation of chemical disorder effects in 24-valence-electron full-Heusler alloys with single magnetic element, Journal of Magnetism and Magnetic Materials 629, 173322 (2025).
- Picozzi et al. [2004] S. Picozzi, A. Continenza, and A. J. Freeman, Role of structural defects on the half-metallic character of and Heusler alloys, Phys. Rev. B 69, 094423 (2004).
- Devi et al. [2018] P. Devi, S. Singh, B. Dutta, K. Manna, S. W. D’Souza, Y. Ikeda, E. Suard, V. Petricek, P. Simon, P. Werner, S. Chadhov, S. S. P. Parkin, C. Felser, and D. Pandey, Adaptive modulation in the magnetic shape-memory Heusler alloy, Phys. Rev. B 97, 224102 (2018).
- Kanomata et al. [2010] T. Kanomata, Y. Chieda, K. Endo, H. Okada, M. Nagasako, K. Kobayashi, R. Kainuma, R. Y. Umetsu, H. Takahashi, Y. Furutani, H. Nishihara, K. Abe, Y. Miura, and M. Shirai, Magnetic properties of the half-metallic Heusler alloys and under pressure, Phys. Rev. B 82, 144415 (2010).
- Suto et al. [2024] H. Suto, V. Barwal, K. Simalaotao, Z. Li, K. Masuda, T. Sasaki, Y. Miura, and Y. Sakuraba, Negative spin polarization of Mn2VGa Heusler alloy thin films studied in current-perpendicular-to-plane giant magnetoresistance devices, Journal of Applied Physics 135, 203901 (2024).
- Ouardi et al. [2013] S. Ouardi, G. H. Fecher, C. Felser, and J. Kübler, Realization of spin gapless semiconductors: The Heusler compound , Phys. Rev. Lett. 110, 100401 (2013).
- Životský et al. [2022] O. Životský, K. Skotnicová, T. Čegan, J. Juřica, L. Gembalová, F. Zažímal, and I. Szurman, Structural and magnetic properties of inverse-Heusler Mn2FeSi alloy powder prepared by ball milling, Materials 15, 697 (2022).
- Wurmehl et al. [2005] S. Wurmehl, G. H. Fecher, H. C. Kandpal, V. Ksenofontov, C. Felser, H.-J. Lin, and J. Morais, Geometric, electronic, and magnetic structure of : Curie temperature and magnetic moment measurements and calculations, Phys. Rev. B 72, 184434 (2005).
- Singh and Mazin [1998] D. J. Singh and I. I. Mazin, Electronic structure, local moments, and transport in , Phys. Rev. B 57, 14352 (1998).
- Khmelevskyi et al. [2015] S. Khmelevskyi, E. Simon, and L. Szunyogh, Antiferromagnetism in full Heusler alloys: Effects of magnetic frustration and chemical disorder, Phys. Rev. B 91, 094432 (2015).
- Sokolovskiy et al. [2012] V. V. Sokolovskiy, V. D. Buchelnikov, M. A. Zagrebin, P. Entel, S. Sahoo, and M. Ogura, First-principles investigation of chemical and structural disorder in magnetic Ni2Mn1+xSn1-x Heusler alloys, Phys. Rev. B 86, 134418 (2012).
- Decolvenaere et al. [2019] E. Decolvenaere, E. Levin, R. Seshadri, and A. Van der Ven, Modeling magnetic evolution and exchange hardening in disordered magnets: The example of Heusler alloys, Phys. Rev. Mater. 3, 104411 (2019).
- Li et al. [2013] T. Li, J. Duan, C. Yang, and X. Kou, Synthesis, microstructure and magnetic properties of Heusler Co2FeSn nanoparticles, Micro & Nano Letters 8, 143 (2013).
- Smith et al. [2024] R. Smith, Z. Gercsi, R. Zhang, K. Siewierska, K. Rode, and J. Coey, Effects of disorder on the magnetic properties of the Heusler alloy V2FeAl, Acta Materialia 267, 119733 (2024).
- Miroshkina et al. [2022] O. N. Miroshkina, B. Eggert, J. Lill, B. Beckmann, D. Koch, M. Y. Hu, T. Lojewski, S. Rauls, F. Scheibel, A. Taubel, M. Šob, K. Ollefs, O. Gutfleisch, H. Wende, M. E. Gruner, and M. Friák, Impact of magnetic and antisite disorder on the vibrational densities of states in Heusler alloys, Phys. Rev. B 106, 214302 (2022).
- Galanakis et al. [2002] I. Galanakis, P. H. Dederichs, and N. Papanikolaou, Slater-Pauling behavior and origin of the half-metallicity of the full-Heusler alloys, Phys. Rev. B 66, 174429 (2002).
- Chico et al. [2016] J. Chico, S. Keshavarz, Y. Kvashnin, M. Pereiro, I. Di Marco, C. Etz, O. Eriksson, A. Bergman, and L. Bergqvist, First-principles studies of the Gilbert damping and exchange interactions for half-metallic Heuslers alloys, Phys. Rev. B 93, 214439 (2016).
- Kuroda et al. [2020] F. Kuroda, T. Fukushima, and T. Oguchi, First-principles study of magnetism and phase stabilities of V2 based antiferromagnetic Heusler alloys, Journal of Applied Physics 127, 193904 (2020).
- Skaftouros et al. [2013] S. Skaftouros, K. Özdoğan, E. Şaşıoğlu, and I. Galanakis, Generalized Slater-Pauling rule for the inverse Heusler compounds, Phys. Rev. B 87, 024420 (2013).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Kresse and Joubert [1999] G. Kresse and D. Joubert, From ultrasoft pseudopotentials to the projector augmented-wave method, Phys. Rev. B 59, 1758 (1999).
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- Hohenberg and Kohn [1964] P. Hohenberg and W. Kohn, Inhomogeneous electron gas, Phys. Rev. 136, B864 (1964).
- Kohn and Sham [1965] W. Kohn and L. J. Sham, Self-consistent equations including exchange and correlation effects, Phys. Rev. 140, A1133 (1965).
- Li et al. [2016] P. Li, X. Liu, M. Chen, P. Lin, X. Ren, L. Lin, C. Yang, and L. He, Large-scale ab initio simulations based on systematically improvable atomic basis, Computational Materials Science 112, 503 (2016), computational Materials Science in China.
- Zhou et al. [2025] W. Zhou, D. Zheng, Q. Liu, D. Lu, Y. Liu, P. Lin, Y. Huang, X. Peng, J. J. Bao, C. Cai, Z. Jin, J. Wu, H. Zhang, G. Jin, Y. Ji, Z. Shen, X. Liu, L. Sun, Y. Cao, M. Sun, J. Liu, T. Chen, R. Liu, Y. Li, H. Han, X. Liang, T. Bao, N. Chen, H. Ren, X. Zhang, Z. Liu, Y. Fu, M. Liu, Z. Li, T. Wen, Z. Tang, Y. Xu, W. Duan, X. Wang, Q. Gu, F.-Z. Dai, Q. Zheng, J. Zhao, Y. Zhang, Q. Ou, H. Jiang, S. Liu, B. Xu, S. Xu, X. Ren, L. He, L. Zhang, and M. Chen, ABACUS: An electronic structure analysis package for the AI era, arXiv e-prints , 2501.08697 (2025).
- Schlipf and Gygi [2015] M. Schlipf and F. Gygi, Optimization algorithm for the generation of ONCV pseudopotentials, Computer Physics Communications 196, 36 (2015).
- He et al. [2021] X. He, N. Helbig, M. J. Verstraete, and E. Bousquet, TB2J: A python package for computing magnetic interaction parameters, Computer Physics Communications 264, 107938 (2021).
- Skubic et al. [2008] B. Skubic, J. Hellsvik, L. Nordström, and O. Eriksson, A method for atomistic spin dynamics simulations: implementation and examples, Journal of Physics: Condensed Matter 20, 315203 (2008).
- Goraus et al. [2023] J. Goraus, M. Fijałkowski, J. Czerniewski, and W. Gumulak, Magnetic properties of V2MnGa, V2MnAl, V2FeGa and V2FeAl, Journal of Magnetism and Magnetic Materials 585, 171103 (2023).
- Zhang et al. [2012] X. Zhang, X. Dai, G. Chen, H. Liu, H. Luo, Y. Li, W. Wang, G. Wu, and G. Liu, First-principle study for full-Heusler compounds V2YAl (Y=V, Cr, Mn, Fe, Co, Ni) and a discussion to Slater–Pauling rule, Computational Materials Science 59, 1 (2012).
- Faleev et al. [2017] S. V. Faleev, Y. Ferrante, J. Jeong, M. G. Samant, B. Jones, and S. S. P. Parkin, Unified explanation of chemical ordering, the Slater-Pauling rule, and half-metallicity in full Heusler compounds, Phys. Rev. B 95, 045140 (2017).
- Xu et al. [2024] Y. Xu, Y.-C. Wang, X. Jin, H. Liu, Y. Liu, H. Song, and F. Tian, Mechanism of magnetic phase transition in correlated magnetic metal: insight into itinerant ferromagnet Fe3‑δGeTe2, Commun. Phys. 7, 381 (2024).
- Xu et al. [2025a] Y. Xu, X. Jin, J. Xiang, H. Zhang, and F. Tian, investigation of the magnetic phase transition in the itinerant ferromagnet , Phys. Rev. B 111, 155142 (2025a).
- Xu et al. [2025b] Y. Xu, X. Jin, H. Tang, and F. Tian, Electronic structures and magnetism in van der waals flat-band material , Europhysics Letters 151, 36002 (2025b).

### v2

@@
-Identifying the magnetic genes in fully- and partially-ordered V2Al ( Cr, Mn, Fe, Co, Ni) Heusler alloys
+Local magnetic structure in fully and partially ordered V2Al Heusler alloys (=Cr, Mn, Fe, Co, Ni)
 Abstract
@@
 In this work, we combine density functional theory with atomistic Monte Carlo simulations to investigate the magnetic ground states, finite-temperature magnetic transitions, and electronic structures of fully-ordered -, -type, and partially-ordered VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys.
-We introduce the concept of magnetic genes, defined as V--V triangular motifs connected by the nearest-neighbor (NN) exchange interactions .
+We propose the concept of magnetic motifs, defined as V--V triangular pathway connected by the nearest-neighbor (NN) exchange interactions .
 Within this framework, the magnetic ground states and transition temperatures across the VAl family can be consistently understood.
-The magnetic order is primarily governed by the NN interactions in the triangular genes, while the transition temperatures are additionally influenced by couplings.
-Furthermore, the magnetic genes are still proven to be effective in our calculations on partially-ordered V2Al alloys from to -type structures.
-Our results suggest that the concept of magnetic genes provides a unifying principle for understanding magnetic ordering in V-based Heusler alloys and could serve as a powerful guide for exploring magnetism and designing advanced spintronic materials in a broader class of Heusler systems.
+The magnetic order is primarily governed by the NN interactions in the triangular motifs, while the transition temperatures are additionally influenced by couplings.
+Furthermore, the magnetic motifs are still proven to be effective in our calculations on partially-ordered V2Al alloys from to -type structures.
+Our results suggest that the concept of magnetic motifs provides a unifying principle for understanding magnetic ordering in V-based Heusler alloys and could serve as a candidate guide for exploring magnetism and designing advanced spintronic materials in a broader class of Heusler systems.
 I INTRODUCTION
@@
 reported that -type V2NbAl and V2TaAl favor antiferromagnetic coupling, whereas in VSi ( Ti, Zr, Hf), the structure often competes with or destabilizes the phase [26].
-In addition, large-scale first-principles screenings of the inverse-Heusler chemical space, where Sc, Ti, V, Cr, Mn; Al, Si, As; and spans from Ti to Zn, have identified the necessary conditions for the emergence of semimetallic states and established three Slater–Pauling rules associated with – hybridization [31].
-Despite these advances, the systematic experimental synthesis and theoretical calculations of V-based VAl compounds with the complete set of Cr, Mn, Fe, Co, and Ni are still lacking.
+Despite these advances, a synthetically theoretical investigation of magentism of V element in VAl compounds remains limited, especially when is considered as a magnetic element.
 A comprehensive microscopic understanding of the mechanisms that govern magnetism in both fully- and partially-ordered V-based full-Heusler alloys is therefore essential.
 In this work, we investigate the ground-state magnetic orders and electronic structures of VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys using density functional theory.
+V atoms play a key role in regulating the magnetic moments of the five magnetic transition elements, influencing both the orientation and magnitude of the magnetic moments.
 For the fully-ordered -type VAl, V2CrAl and V2MnAl exhibit ferrimagnetic order, whereas the remaining compounds adopt ferromagnetic order.
 The situation becomes more intricate in the inverse -type VAl.
-While V2CrAl displays a nearly vanishing net magnetic moment with ferromagnetic order between the two different Wyckoff sites of V atoms, the other four alloys exhibit ferrimagnetic order with antiparallel spin alignment between the two V sublattices.
+While V2CrAl displays a nearly vanishing net magnetic moment with ferromagnetic order between the two V atoms on different Wyckoff sites, the other four alloys exhibit ferrimagnetic order with antiparallel spin alignment between the two V sublattices.
 Despite this diversity, the magnetic orders across VAl compounds can be consistently interpreted in terms of the V--V triangular pathway, governed by the NN magnetic exchange coupling .
@@
 Importantly, we find that the V--V triangular pathway framework remains valid even in partially-ordered systems from to -type structures.
-Building on this insight, we propose the concept of “magnetic genes”, defined as the V--V triangular motif connected by NN exchange interactions .
-We recognize that the magnetic genes could provide a promising concept for exploring the magnetic properties in more Heusler alloys.
+Building on this insight, we propose the concept of “magnetic motifs”, defined as the V--V triangular motif connected by NN exchange interactions .
+We recognize that the magnetic motifs could provide a promising concept for exploring the magnetic properties in more Heusler alloys.
 II METHODS
-To perform the structural relaxation calculations, we use the VASP code [32, 33], which implements the projector-augmented wave (PAW) method [34] and is based on the GGA-PBE generalized gradient approximation [35].
-A dense -mesh of 10×10×10 is employed for the structural relaxation calcuations.
+All calculations were performed based on spin-polarized density functional theory [15, 21].
+To get the equilibrium-state structures of Heusler alloys, we used the VASP code to relax the crystal structures [22, 23], which implements the projector-augmented wave (PAW) method [2], and the generalized gradient approximation in the form of Perdew-Burke-Ernzerhof (GGA-PBE) functional was employed to represent the exchange-correlation potential with a dense -mesh of 10×10×10.
 The plane-wave cutoff was chosen to be 425 eV.
-The convergence criteria for the total energy and ionic forces were set to and , respectively.
-Additionally, we performed spin-polarized density-functional [36, 37] calculations using the Atomic-orbital Based Ab-initio Computation at USTC (ABACUS) package [38, 39], an open-source first-principles platform with advanced ground- and excited-state modules, and interfaces to Wannier90.
-We employed the SG15 optimized norm-conserving Vanderbilt (ONCV) pseudopotentials [40] and the same exchange correlation [35].
-The Kohn-Sham wave functions were expanded using a linear combination of atomic orbitals corresponding to the SG15 ONCV (Version 2.0 standard orbitals) [39].
+The convergence criteria for the total energy and ionic forces were set to and , respectively.For each composition of VAl, we first obtained the relaxed lattice constants of the fully-ordered structure.
+Based on these optimized configurations, partially-ordered structures were constructed by exchanging V and atoms, as illustrated in Fig.
+1(c).
+Each partially ordered model was subsequently subjected to an independent lattice relaxation.
+In this study, all results are obtained from the standard DFT calculations without explicitly accounting for strong electronic correlations.
+Previous DFT investigations have shown that exchange interactions and Curie temperatures in many Heusler alloys can be accurately reproduced without the inclusion of a Hubbard term, yielding good agreement with experimental measurements [24, 40, 35].
+For comparison, the results of DFT+ calculations are provided in the Supplementary Information [1] (see also references [29, 45, 55] therein).
+While the inclusion of the Coulomb interaction does not change the main physical trends reported here, the DFT+ approach generally overestimates the magnitude of the exchange interactions, leading to unrealistically high Curie temperatures.
+We calculated the magnetic orders and electronic structures by using the Atomic-orbital Based Ab-initio Computation at USTC (ABACUS) package [27, 56], an open-source first-principles platform with advanced ground- and excited-state modules, and interfaces to Wannier90.
+We employed the SG15 optimized norm-conserving Vanderbilt (ONCV) pseudopotentials [42] and the same exchange correlation functional GGA-PBE [36].
+The Kohn–Sham electron wave functions were expanded in terms of a linear combination of atomic orbitals (LCAO) associated with the SG15 ONCV pseudopotentials (2.0) [56].
 After convergence tests, the grid cutoff for numerical integrations was set to 100 Ry.
@@
 which was solved with the Vampire atomistic spin-dynamics code [46].
-The simulated system of V2Al is a cubic box of edge length , generated by replicating the relaxed face-centered unit cell and containing about spins, which was evolved with a stochastic spin integrator.
+In Monte Carlo simulations, we examined the convergence with respect to both the supercell size and -point sampling.
+We found that the magnetic transition temperature exhibited negligible variation for supercell sizes ranging from 10 to 20 nm.
+Similarly, for -point grids ranging from to , the results were essentially unchanged for grids finer than .
+Based on these convergence tests, we selected the Monte Carlo supercell of 15 nm and the -point mesh of for all subsequent calculations.
+The simulated system of V2Al is a cubic box of edge length 15 nm, generated by replicating the relaxed face-centered unit cell and containing about spins, which was evolved with a stochastic spin integrator.
 Temperatures from to in increments were sampled using equilibration steps.
@@
 III RESULTS AND DISCUSSION
-III.1 Magnetic orders of fully-ordered - and -V2Al
+III.1 Structural configuration and thermodynamic stability analysis of V2Al
 The magnetic ordering in Heusler alloys is strongly influenced by both their atomic composition and crystallographic structure.
@@
 From a broader perspective, the redistribution of atomic occupations between the 4 and 4 sites, driven by the evolution from to -type structures, highlights the importance of chemical disorder in determining the magnetic properties of these compounds.
+The details of the relaxed lattice parameters for all configurations are summarized in Supplementary Information.
 A detailed analysis of disorder effects will be presented in Sec.
-III.4.
+III.5.
+In the VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys, differences in electronegativity among the constituent atoms play a critical role in determining chemical bonding and, consequently, the structural stability of the compounds.
+The electronegativity difference between vanadium and the -site elements increases from 0.03 to 0.29 Pauli units as varies from Cr to Ni, enhancing the strength of the V- chemical bonds and contributing to improved thermodynamic stability across the series.
+Similarly, the electronegativity difference between Al and the -site atoms rises from 0.05 to 0.31 Pauli units along the same series, providing additional stabilization through -Al chemical bonding.
+In contrast, the electronegativity difference between V and Al is relatively small (0.02 Pauli units), lower than that of both the V- and -Al pairs.
+Nevertheless, this difference still introduces supplementary bonding that help stabilize the overall Heusler structure.
+To assess the stability of the fully ordered and partially ordered V2Al compounds, we calculated both the formation energies and total energies of the possible structural configurations.
+As shown in Fig.
+2(a), all systems, except for the -type V2NiAl, exhibit negative formation energies, indicating their thermodynamic stability.
+These results are consistent with the trends reported in the Open Quantum Materials Database (OQMD) [39, 20].
+The calculated formation energies further reveal that V2MnAl, V2FeAl, and V2CoAl preferentially adopt the -type structure, whereas V2CrAl is most stable in the configuration.
+For V2NiAl, a partially disordered atomic arrangement is energetically favored.
+These conclusions align well with previous theoretical and experimental studies on V-based Heusler alloys [32, 55, 47].
+The corresponding total energy results for both the fully-ordered structures and the partially-ordered structural pathways are presented in Fig.
+2(b).
+Consistent with the formation energy trends, the -type ordering is the lowest-energy configuration for both V2MnAl and V2FeAl.
+Moreover, the total energies of the partially-ordered configurations lie between those of the parent - and -type phases.
+These findings underscore the importance of considering phase stability as a key criterion in the practical design and optimization of high-performance Heusler-based spintronic devices.
+Here, we note that the partially-ordered Heusler alloys are constructed by progressively exchanging V atoms at the 4 sites with atoms at the 4 sites within a 16-atom cubic cell.
+This procedure generates a series of partially-ordered “intermediate” structures.
+However, we find that different disorder configurations for a given composition do not lead to much variation in the exchange interactions.
+Therefore, in this study, we present the results for the configuration with the lowest total energy.
 | Compound | Source | |||||
@@
 | -V2NiAl | -0.66 | 1.74 | 0.06 | -0.06 | 1.08 | Zhang et al. [55] |
+III.2 Magnetic orders of fully-ordered - and -V2Al
 We begin by examining the effect of atomic composition on the magnetic order in -type VAl ( Cr, Mn, Fe, Co, Ni), where the crystal structure is relatively simple, containing only three inequivalent Wyckoff sites compared to the more complex -type structure.
-As summarized in Table 1, compounds with Fe, Co, and Ni exhibit ferromagnetic ordering, whereas those with Cr and Mn display ferrimagnetic ordering characterized by antiparallel spin alignment between the and V sublattices.
+As summarized in Table 1, compounds with Fe, Co and Ni exhibit ferromagnetic ordering, whereas those with Cr and Mn display ferrimagnetic ordering characterized by antiparallel spin alignment between the and V sublattices.
 In particular, our calculations show that V2MnAl adopts a ferrimagnetic ground state, with magnetic moments of per V atom and per Mn atom.
@@
 A detailed discussion of the interplay between electronic structure and magnetic order will be presented in Sec.
-III.5.
+III.6.
 To investigate the influence of lattice type on magnetism, we next examine the -type inverse-Heusler VAl compounds, which possess a more complex crystal structure with four inequivalent Wyckoff sites compared to the -type structure.
@@
 These results highlight the intricate and composition-dependent magnetic behaviors of inverse-Heusler VAl compounds, motivating the need for a more general understanding of their underlying mechanisms.
-III.2 Microscopic magnetic mechanism in - and -V2Al
+III.3 Microscopic magnetic mechanism in - and -V2Al
 To gain deeper insight into the origin of magnetism in these systems, we investigate the magnetic exchange couplings () across different atomic compositions and lattice types.
@@
 Correspondingly, as shown in Figs.
-2(a) and 2(b), the NN exchange coupling () between Mn (Cr) atoms at the 4 site and V atoms at the 8 site is antiferromagnetic type, which has the largest value among all the other magnetic exchange couplings in these materials.
-In contrast, for Fe, Co, and Ni, the NN coupling between the and V atoms is ferromagnetic, as illustrated in Fig.
-2(c).
+3(a) and 3(b), the NN exchange coupling () between Mn (Cr) atoms at the 4 site and V atoms at the 8 site is antiferromagnetic type, which has the largest value among all the other magnetic exchange couplings in these materials.
+In contrast, for Fe, Co and Ni, the NN coupling between the and V atoms is ferromagnetic, as illustrated in Fig.
+3(c).
 Meanwhile, the exchange interactions between - atoms are significantly weaker and do not play a decisive role in stabilizing the magnetic order.
@@
 In V2MnAl, both the NN and next-nearest-neighbor (NNN) V-V couplings, indicated by the gray line in Fig.
-2(a), are ferromagnetic.
+3(a), are ferromagnetic.
 However, the NN V-V coupling in V2CrAl, colored by the gray line, acquires a finite antiferromagnetic contribution, as shown in Fig.
-2(b).
+3(b).
 Taken together, these results indicate that antiferromagnetic V-Mn(Cr)-V exchange pathways form triangular motifs that ultimately stabilize ferromagnetic ordering among the V sublattice.
@@
 As shown in Figs.
-2(d) and 2(e), the NN exchange couplings (highlighted in blue) are strongly antiferromagnetic ( Cr, Mn; couplings for Fe, Co, Ni are omitted for clarity) and represent the primary factor stabilizing ferrimagnetic ordering between the and V1 sublattices.
+3(d) and 3(e), the NN exchange couplings (highlighted in blue) are strongly antiferromagnetic ( Cr, Mn; couplings for Fe, Co, Ni are shown in Supplementary Information) and represent the primary factor stabilizing ferrimagnetic ordering between the and V1 sublattices.
 In contrast, the NN exchange coupling between and V2 (4 site), as shown in purple lines in Figs.
-2(e) and 2(f), evolves rapidly from antiferromagnetic to ferromagnetic as changes from Cr to Ni.
+3(e) and 3(f), evolves rapidly from antiferromagnetic to ferromagnetic as changes from Cr to Ni.
 This change of is consistent with the corresponding magnetic ordering from ferrimagnetic for Cr to ferromagnetic for Mn, Fe, Co, and Ni.
@@
 In -type VAl, the diverse magnetic orders can be rationalized in terms of triangular exchange pathways formed by V--V atoms, as illustrated in the inset of Fig.
-2(a).
+3(a).
 In -type VAl, the underlying mechanism is similar, but an additional type of magnetic exchange coupling is present.
 Our results indicate that the V1--V2 triangular pathway remains a key determinant of magnetic order in -type compounds, as shown in the inset of Fig.
-2(d).
+3(d).
 Changes in the type of the coupling substantially influence the magnetic ordering in these materials.
 Moreover, although the direct NN exchange coupling between - atoms ( Mn, Fe, Co, Ni) is antiferromagnetic type in -VAl, the triangular -V1- pathway ultimately stabilizes ferromagnetic ordering due to the intervening antiferromagnetic X-V1 couplings, as illustrated in the inset of Fig.
-2(d).
+3(d).
 In summary, despite variations in atomic composition and lattice type, the site plays a central role in controlling the magnetic properties of V-based VAl compounds through triangular pathways formed by NN V and atoms.
-We therefore identify these motif as a candidate “magnetic genes”.
-III.3 Magnetic transition temperature of - and -V2Al
+We therefore identify these pathway as a candidate “magnetic motifs”.
+III.4 Magnetic transition temperature of - and -V2Al
 The microscopic mechanisms of magnetic exchange couplings directly determine the macroscopic magnetic properties.
@@
 As shown in Fig.
-3(a), in -type VAl, V2MnAl exhibits the highest Curie temperature, reaching approximately 900K, whereas the Curie temperatures of the remaining compounds ( Cr, Fe, Co, Ni) are significantly lower, ranging from 100 K to 300 K.
+4(a), in -type VAl, V2MnAl exhibits the highest Curie temperature, reaching approximately 900 K, whereas the Curie temperatures of the remaining compounds ( Cr, Fe, Co, Ni) are significantly lower, ranging from 100 K to 300 K.
 The exceptionally high Curie temperature of V2MnAl originates from the effective ferromagnetic coupling between V atoms mediated by the triangular V-Mn-V pathways, which involve strong antiferromagnetic NN couplings ( meV).
@@
 Specifically, as shown in Fig.
-3(a), decreases from 900 K for V2MnAl to 100 K for V2FeAl, consistent with the corresponding NN exchange couplings , which vary from to 2.0 meV, as illustrated in Fig.
-2.
+4(a), decreases from 900 K for V2MnAl to 100 K for V2FeAl, consistent with the corresponding NN exchange couplings , which vary from to 2.0 meV, as illustrated in Fig.
+3.
 The temperature dependence of magnetic phase transitions in -type VAl compounds differs markedly from that in -type structures.
@@
 In V2CrAl, the ferromagnetic – interactions enhance , whereas in V2MnAl, the coupling is antiferromagnetic, leading to a notable reduction of .
-In the series Fe, Co, Ni, the gradual decrease of is primarily driven by the weakening of the NN coupling, which diminishes from to 0 meV.
+In the series Fe, Co, Ni, the gradual decrease of is primarily driven by the weakening the absolute value of the NN coupling, which diminishes from to 0.1 meV (see Supplementary Information).
 In summary, the magnetic ground states and magnetic transition temperatures of VAl compounds can be rationalized within the framework of the V--V triangular exchange pathway and its associated NN magnetic exchange couplings, applicable to both - and -type structures.
@@
 | 100% | ––– | 1.610 | -0.820 | ––– | 1.466 | 0.007 | 2.264 |
-III.4 Effect of chemical disorder on magnetism of partially-ordered V2Al
+III.5 Effect of chemical disorder on magnetism of partially-ordered V2Al
 In VAl compounds, the NN magnetic exchange coupling plays a central role in determining both the magnetic ground states and the magnetic transition temperatures.
@@
 As shown in Fig.
-4(a), the Curie temperature decreases from 900 K in the fully-ordered structure to 420 K in the configuration.
+5(a), the Curie temperature decreases from 900 K in the fully-ordered structure to 420 K in the configuration.
 Within our framework, the NN magnetic exchange couplings and are the primary factors governing the magnetic transition temperature.
@@
 As shown in Fig.
-5(a), the Curie temperature of V2FeAl gradually decreases from 600 K to 120 K as the structure evolves from - to type.
+6(a), the Curie temperature of V2FeAl gradually decreases from 600 K to 120 K as the structure evolves from - to type.
 This observation raises the question of which factors primarily govern this anomalous sequence of magnetic transition temperatures.
@@
 As shown in Fig.
-5(b), the absolute value of decreases systematically from the - to the -type structure, consistent with the observed reduction in .
+6(b), the absolute value of decreases systematically from the - to the -type structure, consistent with the observed reduction in .
 These results confirm the validity of as the primary determinant of the Curie temperature in partially-ordered V2FeAl.
-Our calculations on partially chemically disordered VAl indicate that the NN magnetic exchange coupling remains the dominant factor governing the magnetic transition temperature.
+Our calculations on partially chemically disordered VAl (for Cr, Co, Ni, see Supplementary Information) indicate that the NN magnetic exchange coupling remains the dominant factor governing the magnetic transition temperature.
 Notably, during the transition from the to the -type structure, the 4 site is consistently occupied by V2 atoms, ensuring that the V-X-V magnetic triangular pathway is preserved.
 Consequently, the microscopic magnetic mechanism in chemically disordered VAl is still primarily governed by the V-X-V triangular pathway and the associated NN magnetic exchange couplings .
-III.5 Electronic structures of fully-ordered V2Al
+III.6 Electronic structures of fully-ordered V2Al
 Based on the calculated magnetic ground states of -type VAl in Sec.
-III.1, the observed complex magnetic orders and non-monotonic variation of total magnetic moments with atomic composition are proposed related to their electronic structures.
+III.2, the observed complex magnetic orders and non-monotonic variation of total magnetic moments with atomic composition are proposed related to their electronic structures.
 In -type VAl, the relatively small total magnetic moments of V2CrAl and V2MnAl originate from the anti-parallel alignment of spin moments between the V and sublattices.
@@
 As shown in Fig.
-6(a), the total DOS of -type V2Al exhibits a high value at the Fermi level, indicating itinerant magnetism in these compounds [46, 47, 48].
+7(a), the total DOS of -type V2Al exhibits a high value at the Fermi level, indicating itinerant magnetism in these compounds [54, 53, 52].
 Notably, several compounds display nearly half-metallic character.
 As shown in Fig.
-6(b), Cr and Mn atoms exhibit pronounced occupied states with sharp peaks at approximately -1 eV and -2 eV in the spin-down channel, while their spin-up channel states are minimally occupied.
+7(b), Cr and Mn atoms exhibit pronounced occupied states with sharp peaks at approximately -1 eV and -2 eV in the spin-down channel, while their spin-up channel states are minimally occupied.
 In contrast, for Fe, Co, and Ni, the PDOS of atoms shows significant occupation in the spin-up channel and comparatively low values in the spin-down channel.
@@
 Furthermore, from Fe to Ni, the PDOS of atoms shifts to lower energies in both spin-up and spin-down channels, as shown in Fig.
-6(b).
+7(b).
 This trend is attributed to the increasing number of valence electrons with rising atomic number.
@@
 Examination of the V-atom PDOS from Fe to Ni reveals an increase in spin-up states below the Fermi level for V atoms in Co- and Ni-containing compounds, particularly in the energy range from -1 to 0 eV, as shown in Fig.
-6(c), while the spin-down PDOS decreases in the same energy range.
+7(c), while the spin-down PDOS decreases in the same energy range.
 This behavior correlates with the observed increase in magnetic moments of V atoms from Fe to Ni.
 In contrast, the PDOS of Al atoms remains largely unchanged across the series in Fig.
-6(d), consistent with the negligible magnetic moment on Al atoms.
+7(d), consistent with the negligible magnetic moment on Al atoms.
 Taken together, the change in magnetic order from Mn to Fe, combined with the opposing trends in magnetic moments of V and atoms from Fe to Ni, accounts for the non-monotonic variation in the total magnetic moment across the V2Al series.
+From our calculated electronic structures, we observe substantial hybridization among the V-, -, and Al- states near the Fermi level in all compounds, as evidenced by the similar peak shapes and their alignment in energy.
+Although such hybridization is an evident feature throughout the VAl series, the key factor governing the evolution of magnetic behavior is the variation in band filling associated with the -site element.
+A comparison of the DOS for Cr and Mn with those for Fe, Co, and Ni clearly shows that, in the Cr and Mn systems, the -electron manifold lies much closer to half-filling.
+This near half-filled configuration enhances antiferromagnetic superexchange between V and Cr (Mn) atoms through virtual electron-hopping processes between neighboring sites [41, 38].
+In contrast, as the -site element becomes more electron-rich (from Fe to Ni), the increased band filling shifts the DOS away from half-filling, thereby suppressing the antiferromagnetic superexchange channel and promoting ferromagnetic exchange.
 IV CONCLUSION
-In this work, we employed density functional theory calculations combined with Monte Carlo simulations based on Heisenberg magnetic exchange couplings to investigate V2Al alloys and introduced the concept of “magnetic genes”.
-These magnetic genes govern both the magnetic ground states and the magnetic transition temperatures in fully- and partially-ordered compounds.
+In this work, we employed density functional theory calculations combined with Monte Carlo simulations based on Heisenberg magnetic exchange couplings to investigate V2Al alloys and introduced the concept of “magnetic motifs”.
+These magnetic motifs govern both the magnetic ground states and the magnetic transition temperatures in fully- and partially-ordered compounds.
 The underlying physical mechanism arises from the modulation of NN magnetic exchange couplings within the V--V triangular pathways, which is influenced by atomic composition, crystal structure, and the degree of chemical disorder.
-Using the framework of magnetic genes, the magnetic orders can be rationalized in terms of the V--V triangular pathways formed by NN couplings.
+Using the framework of magnetic motifs, the magnetic orders can be rationalized in terms of the V--V triangular pathways formed by NN couplings.
 Furthermore, the magnetic transition temperatures are predominantly determined by both NN and additional magnetic exchange interactions, irrespective of whether the alloy is fully or partially ordered.
-We anticipate that the identification of magnetic genes in other Heusler alloys may provide a systematic and predictive approach to understanding and tuning their diverse magnetic properties.
+The concept of “magnetic motifs” captures the minimal set of dominant exchange interactions that governs both the magnetic ground states and the evolution of the Curie temperature across different structures and compositions.
+We anticipate that identifying such magnetic motifs in other Heusler alloys will provide a systematic and predictive framework for understanding and tailoring their diverse magnetic properties.
+Furthermore, a comprehensive investigation that combines phase stability analysis with detailed magnetic characterization will be essential for guiding the design of high-performance Heusler-based spintronic devices in the future studies.
 Acknowledgements.
@@
 52371174 and No.
-12204033), the Fundamental Research Funds for the Central Universities (Grant No.
-FRF-TP-22-097A1), the State Key Lab of Advanced Metals and Materials (Grant No.
-2022Z-13), the Young Elite Scientist Sponsorship Program by BAST (Grant No.
-BYESS2023301) and the High-performance Computing Platform of University of Science and Technology Beijing.References
-- Gupta et al.
-[2023] S.
-Gupta, S.
-Chakraborty, V.
-Bhasin, S.
-Pakhira, S.
-Dan, C.
-Barreteau, J.-C.
-Crivello, S.
-N.
-Jha, M.
-Avdeev, J.-M.
-Greneche, D.
-Bhattacharyya, E.
-Alleno, and C.
-Mazumdar, High spin polarization in the disordered quaternary heusler alloy femnvga, Phys.
+12204033), Science Challenge Project (Grant No.TZ2025009), the State Key Lab of Advanced Metals and Materials (Grant No.
+2025Z-Z22).
+Numerical computations were performed at the High-performance Computing Platform of University of Science and Technology Beijing.References
+- [1] Note: See Supplemental Material at [URL will be inserted by publisher] for more theoretical analysis. Cited by: §II.
+- [2] (1994-12) Projector augmented-wave method. Phys. Rev. B 50, pp. 17953–17979. External Links: Document, Link Cited by: §II.
+- [3] (2016-06) First-principles studies of the Gilbert damping and exchange interactions for half-metallic Heuslers alloys. Phys. Rev. B 93, pp. 214439. External Links: Document, Link Cited by: §I.
+- [4] (1983-06) New class of materials: half-metallic ferromagnets. Phys. Rev. Lett. 50, pp. 2024–2027. External Links: Document, Link Cited by: §I.
+- [5] (2019-10) Modeling magnetic evolution and exchange hardening in disordered magnets: the example of Heusler alloys. Phys. Rev. Mater. 3, pp. 104411. External Links: Document, Link Cited by: §I.
+- [6] (2018-06) Adaptive modulation in the magnetic shape-memory Heusler alloy. Phys. Rev. B 97, pp. 224102. External Links: Document, Link Cited by: §I.
+- [7] (2017-01) Unified explanation of chemical ordering, the Slater-Pauling rule, and half-metallicity in full Heusler compounds.
+Phys.
 Rev.
-B 108, 045137 (2023).
-- Venkateswara et al.
-[2023] Y.
-Venkateswara, J.
-Nag, S.
-S.
-Samatham, A.
-K.
-Patel, P.
-D.
-Babu, M.
-R.
-Varma, J.
-Nayak, K.
-G.
-Suresh, and A.
-Alam, FeRhCrSi: Spin semimetal with spin-valve behavior at room temperature, Phys.
+B 95, pp.
+045140.
+External Links: Document, Link Cited by: §III.2.
+- [8] (2024) High-throughput screening of all--metal Heusler alloys for magnetocaloric applications. Chem. Mater. 36 (14), pp. 6765–6776. External Links: Document, Link Cited by: §I.
+- [9] (2002-11) Slater-Pauling behavior and origin of the half-metallicity of the full-Heusler alloys. Phys. Rev. B 66, pp. 174429. External Links: Document, Link Cited by: §I, §III.1.
+- [10] (2023) Magnetic properties of V2MnGa, V2MnAl, V2FeGa and V2FeAl.
+J.
+Magn.
+Magn.
+Mater.
+585, pp.
+171103.
+External Links: ISSN 0304-8853, Document, Link Cited by: Table 1, Table 1, Table 1, §III.2, §III.2, Table 2, Table 2.
+- [11] (2011) Simple rules for the understanding of Heusler compounds. Prog. Solid State Chem. 39 (1), pp. 1–50. External Links: ISSN 0079-6786, Document, Link Cited by: §I, §III.1.
+- [12] (2025) Machine learning investigation of chemical disorder effects in 24-valence-electron full-Heusler alloys with single magnetic element.
+J.
+Magn.
+Magn.
+Mater.
+629, pp.
+173322.
+External Links: ISSN 0304-8853, Document, Link Cited by: §I.
+- [13] (2023-07) High spin polarization in the disordered quaternary heusler alloy femnvga. Phys. Rev. B 108, pp. 045137. External Links: Document, Link Cited by: §I.
+- [14] (2021) TB2J: a python package for computing magnetic interaction parameters. Comput. Phys. Commun. 264, pp. 107938. External Links: ISSN 0010-4655, Document, Link Cited by: §II.
+- [15] (1964-11) Inhomogeneous electron gas. Phys. Rev. 136, pp. B864–B871. External Links: Document, Link Cited by: §II.
+- [16] (2006-05) Temperature dependence of the anomalous hall conductivity in the Heusler alloy . Phys. Rev. B 73, pp. 172417. External Links: Document, Link Cited by: §I.
+- [17] (2010-10) Magnetic properties of the half-metallic Heusler alloys and under pressure. Phys. Rev. B 82, pp. 144415. External Links: Document, Link Cited by: §I.
+- [18] (2008-04) Half-metallic ferromagnets: from band structure to many-body effects. Rev. Mod. Phys. 80, pp. 315–378. External Links: Document, Link Cited by: §I.
+- [19] (2015-03) Antiferromagnetism in full Heusler alloys: effects of magnetic frustration and chemical disorder. Phys. Rev. B 91, pp. 094432. External Links: Document, Link Cited by: §I.
+- [20] (2015-12) The Open Quantum Materials Database (OQMD): assessing the accuracy of DFT formation energies.
+npj Computational Materials 1 (1), pp.
+15010.
+External Links: Document, Link Cited by: §III.1.
+- [21] (1965-11) Self-consistent equations including exchange and correlation effects. Phys. Rev. 140, pp. A1133–A1138. External Links: Document, Link Cited by: §II.
+- [22] (1996-10) Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set. Phys. Rev. B 54, pp. 11169–11186. External Links: Document, Link Cited by: §II.
+- [23] (1999-01) From ultrasoft pseudopotentials to the projector augmented-wave method. Phys. Rev. B 59, pp. 1758–1775. External Links: Document, Link Cited by: §II.
+- [24] (2007-07) Understanding the trend in the Curie temperatures of -based Heusler compounds: Ab initio calculations. Phys. Rev. B 76, pp. 024414. External Links: Document, Link Cited by: §II.
+- [25] (2007-07) Understanding the trend in the curie temperatures of -based Heusler compounds: ab initio calculations. Phys. Rev. B 76, pp. 024414. External Links: Document, Link Cited by: §I.
+- [26] (2020-05) First-principles study of magnetism and phase stabilities of V2 based antiferromagnetic Heusler alloys.
+J.
+Appl.
+Phys.
+127 (19), pp.
+193904.
+External Links: ISSN 0021-8979, Document, Link Cited by: §I.
+- [27] (2016) Large-scale ab initio simulations based on systematically improvable atomic basis.
+Comput.
+Mater.
+Sci.
+112, pp.
+503–517.
+Note: Computational Materials Science in China External Links: ISSN 0927-0256, Document, Link Cited by: §II.
+- [28] (2013) Synthesis, microstructure and magnetic properties of Heusler Co2FeSn nanoparticles. Micro & Nano Lett. 8 (3), pp. 143–146. External Links: Document, Link Cited by: §I.
+- [29] (2023-10) Doubly Screened Coulomb Correction approach for strongly correlated systems. J. Phys. Chem. Lett. 14, pp. 8930–8939. External Links: Document, Link Cited by: §II.
+- [30] (2023-05) Exploration of all- Heusler alloys for permanent magnets: an ab initio based high-throughput study. Phys. Rev. B 107, pp. 174402. External Links: Document, Link Cited by: §I.
+- [31] (2022-12) Impact of magnetic and antisite disorder on the vibrational densities of states in Heusler alloys. Phys. Rev. B 106, pp. 214302. External Links: Document, Link Cited by: §I.
+- [32] (2024) Orbital-engineered anomalous Hall conductivity in stable full Heusler compounds: a pathway to optimized spintronics.
+Comput.
+Mater.
+Sci.
+12 (37), pp.
+15082–15091.
+External Links: Document Cited by: §III.1.
+- [33] (2025-04) Near-half-metallic state in the half-Heusler PtMnSb film on a III-V substrate. Phys. Rev. Mater. 9, pp. 044403. External Links: Document, Link Cited by: §I.
+- [34] (2013-03) Realization of spin gapless semiconductors: the Heusler compound . Phys. Rev. Lett. 110, pp. 100401. External Links: Document, Link Cited by: §I.
+- [35] (2001-10) Ab initio calculations of exchange interactions, spin-wave stiffness constants, and Curie temperatures of Fe, Co, and Ni.
+Phys.
 Rev.
-B 107, L100401 (2023).
-- Nishihaya et al.
-[2025] S.
-Nishihaya, M.
+B 64, pp.
+174402.
+External Links: Document, Link Cited by: §II.
+- [36] (1996-10) Generalized gradient approximation made simple. Phys. Rev. Lett. 77, pp. 3865–3868. External Links: Document, Link Cited by: §II.
+- [37] (2004-03) Role of structural defects on the half-metallic character of and Heusler alloys. Phys. Rev. B 69, pp. 094423. External Links: Document, Link Cited by: §I, §III.1.
+- [38] (2013-04) Antiferromagnetic exchange interactions in the Ni2Mn1.4In0.6 ferromagnetic Heusler alloy. Phys. Rev. B 87, pp. 144412. External Links: Document, Link Cited by: §III.6.
+- [39] (2013-11) Materials design and discovery with high-throughput density functional theory: the Open Quantum Materials Database (OQMD).
+JOM 65 (11), pp.
+1501–1509.
+External Links: Document, Link Cited by: §III.1.
+- [40] (2005-11) Exchange interactions and temperature dependence of magnetization in half-metallic Heusler alloys. Phys. Rev. B 72, pp. 184415. External Links: Document, Link Cited by: §II.
+- [41] (2008-02) Role of conduction electrons in mediating exchange interactions in Mn-based Heusler alloys. Phys. Rev. B 77, pp. 064417. External Links: Document, Link Cited by: §III.6.
+- [42] (2015) Optimization algorithm for the generation of ONCV pseudopotentials. Comput. Phys. Commun. 196, pp. 36–44. External Links: ISSN 0010-4655, Document, Link Cited by: §II.
+- [43] (2018-03) Pressure effect on the magnetic properties of the half-metallic Heusler alloy . Phys. Rev. B 97, pp. 104414. External Links: Document, Link Cited by: §I.
+- [44] (1998-06) Electronic structure, local moments, and transport in . Phys. Rev. B 57, pp. 14352–14356. External Links: Document, Link Cited by: §I.
+- [45] (2013-01) Generalized Slater-Pauling rule for the inverse Heusler compounds. Phys. Rev. B 87, pp. 024420. External Links: Document, Link Cited by: §II, §III.2, Table 2.
+- [46] (2008-07) A method for atomistic spin dynamics simulations: implementation and examples. J. Phys.: Condens. Matter 20 (31), pp. 315203. External Links: Document, Link Cited by: §II.
+- [47] (2024) Effects of disorder on the magnetic properties of the Heusler alloy V2FeAl. Acta Mater. 267, pp. 119733. External Links: ISSN 1359-6454, Document, Link Cited by: §I, §III.1.
+- [48] (2012-10) First-principles investigation of chemical and structural disorder in magnetic Ni2Mn1+xSn1-x Heusler alloys. Phys. Rev. B 86, pp. 134418. External Links: Document, Link Cited by: §I.
+- [49] (2024-05) Negative spin polarization of Mn2VGa Heusler alloy thin films studied in current-perpendicular-to-plane giant magnetoresistance devices.
 J.
-A.
-Jardine, H.
-S.
-Inbar, A.
-Goswami, J.
-T.
-Dong, A.
-N.
-Engel, Y.-H.
-Chang, C.
-P.
-Dempsey, M.
-Hashimoto, D.
-Lu, N.
-Marom, and C.
-J.
-Palmstrøm, Near-half-metallic state in the half-Heusler PtMnSb film on a III-V substrate, Phys.
+Appl.
+Phys.
+135 (20), pp.
+203901.
+External Links: ISSN 0021-8979, Document, Link Cited by: §I.
+- [50] (2023-03) FeRhCrSi: spin semimetal with spin-valve behavior at room temperature. Phys. Rev. B 107, pp. L100401. External Links: Document, Link Cited by: §I.
+- [51] (2005-11) Geometric, electronic, and magnetic structure of : curie temperature and magnetic moment measurements and calculations.
+Phys.
 Rev.
+B 72, pp.
+184434.
+External Links: Document, Link Cited by: §I.
+- [52] (2025-08) Electronic structures and magnetism in van der waals flat-band material . Europhys. Lett. 151 (3), pp. 36002. External Links: Document, Link Cited by: §III.6.
+- [53] (2025-04) Investigation of the magnetic phase transition in the itinerant ferromagnet . Phys. Rev. B 111, pp. 155142. External Links: Document, Link Cited by: §III.6.
+- [54] (2024-11) Mechanism of magnetic phase transition in correlated magnetic metal: insight into itinerant ferromagnet Fe3‑δGeTe2.
+Commun.
+Phys.
+7 (1), pp.
+381.
+External Links: Document Cited by: §III.6.
+- [55] (2012) First-principle study for full-Heusler compounds V2YAl (Y=V, Cr, Mn, Fe, Co, Ni) and a discussion to Slater–Pauling rule.
+Comput.
 Mater.
-9, 044403 (2025).
-- de Groot et al. [1983] R. A. de Groot, F. M. Mueller, P. G. v. Engen, and K. H. J. Buschow, New class of materials: Half-metallic ferromagnets, Phys. Rev. Lett. 50, 2024 (1983).
-- Žutić et al. [2004] I. Žutić, J. Fabian, and S. Das Sarma, Spintronics: Fundamentals and applications, Rev. Mod. Phys. 76, 323 (2004).
-- Katsnelson et al.
-[2008] M.
-I.
-Katsnelson, V.
-Y.
-Irkhin, L.
-Chioncel, A.
-I.
-Lichtenstein, and R.
-A.
-de Groot, Half-metallic ferromagnets: From band structure to many-body effects, Rev.
-Mod.
-Phys.
-80, 315 (2008).
-- Graf et al. [2011] T. Graf, C. Felser, and S. S. Parkin, Simple rules for the understanding of Heusler compounds, Progress in Solid State Chemistry 39, 1 (2011).
-- Husmann and Singh [2006] A. Husmann and L. J. Singh, Temperature dependence of the anomalous hall conductivity in the Heusler alloy , Phys. Rev. B 73, 172417 (2006).
-- Shigeta et al.
-[2018] I.
-Shigeta, Y.
-Fujimoto, R.
-Ooka, Y.
-Nishisako, M.
-Tsujikawa, R.
-Y.
-Umetsu, A.
-Nomura, K.
-Yubuta, Y.
-Miura, T.
-Kanomata, M.
-Shirai, J.
-Gouchi, Y.
-Uwatoko, and M.
-Hiroi, Pressure effect on the magnetic properties of the half-metallic Heusler alloy , Phys.
-Rev.
-B 97, 104414 (2018).
-- Kübler et al. [2007] J. Kübler, G. H. Fecher, and C. Felser, Understanding the trend in the curie temperatures of -based Heusler compounds: Ab initio calculations, Phys. Rev. B 76, 024414 (2007).
-- Fortunato et al.
-[2024] N.
-M.
-Fortunato, X.
-Li, S.
-Schönecker, R.
-Xie, A.
-Taubel, F.
-Scheibel, I.
-Opahle, O.
-Gutfleisch, and H.
-Zhang, High-throughput screening of all--metal Heusler alloys for magnetocaloric applications, Chemistry of Materials 36, 6765 (2024).
-- Marathe and Herper [2023] M. Marathe and H. C. Herper, Exploration of all- Heusler alloys for permanent magnets: An ab initio based high-throughput study, Phys. Rev. B 107, 174402 (2023).
-- Guo et al.
-[2025] Y.
-Guo, C.
-Guo, Y.
-Xu, B.
-Zheng, X.
-Ni, and F.
-Tian, Machine learning investigation of chemical disorder effects in 24-valence-electron full-Heusler alloys with single magnetic element, Journal of Magnetism and Magnetic Materials 629, 173322 (2025).
-- Picozzi et al. [2004] S. Picozzi, A. Continenza, and A. J. Freeman, Role of structural defects on the half-metallic character of and Heusler alloys, Phys. Rev. B 69, 094423 (2004).
-- Devi et al.
-[2018] P.
-Devi, S.
-Singh, B.
-Dutta, K.
-Manna, S.
-W.
-D’Souza, Y.
-Ikeda, E.
-Suard, V.
-Petricek, P.
-Simon, P.
-Werner, S.
-Chadhov, S.
-S.
-P.
-Parkin, C.
-Felser, and D.
-Pandey, Adaptive modulation in the magnetic shape-memory Heusler alloy, Phys.
-Rev.
-B 97, 224102 (2018).
-- Kanomata et al.
-[2010] T.
-Kanomata, Y.
-Chieda, K.
-Endo, H.
-Okada, M.
-Nagasako, K.
-Kobayashi, R.
-Kainuma, R.
-Y.
-Umetsu, H.
-Takahashi, Y.
-Furutani, H.
-Nishihara, K.
-Abe, Y.
-Miura, and M.
-Shirai, Magnetic properties of the half-metallic Heusler alloys and under pressure, Phys.
-Rev.
-B 82, 144415 (2010).
-- Suto et al.
-[2024] H.
-Suto, V.
-Barwal, K.
-Simalaotao, Z.
-Li, K.
-Masuda, T.
-Sasaki, Y.
-Miura, and Y.
-Sakuraba, Negative spin polarization of Mn2VGa Heusler alloy thin films studied in current-perpendicular-to-plane giant magnetoresistance devices, Journal of Applied Physics 135, 203901 (2024).
-- Ouardi et al. [2013] S. Ouardi, G. H. Fecher, C. Felser, and J. Kübler, Realization of spin gapless semiconductors: The Heusler compound , Phys. Rev. Lett. 110, 100401 (2013).
-- Životský et al.
-[2022] O.
-Životský, K.
-Skotnicová, T.
-Čegan, J.
-Juřica, L.
-Gembalová, F.
-Zažímal, and I.
-Szurman, Structural and magnetic properties of inverse-Heusler Mn2FeSi alloy powder prepared by ball milling, Materials 15, 697 (2022).
-- Wurmehl et al.
-[2005] S.
-Wurmehl, G.
-H.
-Fecher, H.
-C.
-Kandpal, V.
-Ksenofontov, C.
-Felser, H.-J.
-Lin, and J.
-Morais, Geometric, electronic, and magnetic structure of : Curie temperature and magnetic moment measurements and calculations, Phys.
-Rev.
-B 72, 184434 (2005).
-- Singh and Mazin [1998] D. J. Singh and I. I. Mazin, Electronic structure, local moments, and transport in , Phys. Rev. B 57, 14352 (1998).
-- Khmelevskyi et al. [2015] S. Khmelevskyi, E. Simon, and L. Szunyogh, Antiferromagnetism in full Heusler alloys: Effects of magnetic frustration and chemical disorder, Phys. Rev. B 91, 094432 (2015).
-- Sokolovskiy et al.
-[2012] V.
-V.
-Sokolovskiy, V.
-D.
-Buchelnikov, M.
-A.
-Zagrebin, P.
-Entel, S.
-Sahoo, and M.
-Ogura, First-principles investigation of chemical and structural disorder in magnetic Ni2Mn1+xSn1-x Heusler alloys, Phys.
-Rev.
-B 86, 134418 (2012).
-- Decolvenaere et al.
-[2019] E.
-Decolvenaere, E.
-Levin, R.
-Seshadri, and A.
-Van der Ven, Modeling magnetic evolution and exchange hardening in disordered magnets: The example of Heusler alloys, Phys.
-Rev.
-Mater.
-3, 104411 (2019).
-- Li et al. [2013] T. Li, J. Duan, C. Yang, and X. Kou, Synthesis, microstructure and magnetic properties of Heusler Co2FeSn nanoparticles, Micro & Nano Letters 8, 143 (2013).
-- Smith et al.
-[2024] R.
-Smith, Z.
-Gercsi, R.
-Zhang, K.
-Siewierska, K.
-Rode, and J.
-Coey, Effects of disorder on the magnetic properties of the Heusler alloy V2FeAl, Acta Materialia 267, 119733 (2024).
-- Miroshkina et al.
-[2022] O.
-N.
-Miroshkina, B.
-Eggert, J.
-Lill, B.
-Beckmann, D.
-Koch, M.
-Y.
-Hu, T.
-Lojewski, S.
-Rauls, F.
-Scheibel, A.
-Taubel, M.
-Šob, K.
-Ollefs, O.
-Gutfleisch, H.
-Wende, M.
-E.
-Gruner, and M.
-Friák, Impact of magnetic and antisite disorder on the vibrational densities of states in Heusler alloys, Phys.
-Rev.
-B 106, 214302 (2022).
-- Galanakis et al. [2002] I. Galanakis, P. H. Dederichs, and N. Papanikolaou, Slater-Pauling behavior and origin of the half-metallicity of the full-Heusler alloys, Phys. Rev. B 66, 174429 (2002).
-- Chico et al.
-[2016] J.
-Chico, S.
-Keshavarz, Y.
-Kvashnin, M.
-Pereiro, I.
-Di Marco, C.
-Etz, O.
-Eriksson, A.
-Bergman, and L.
-Bergqvist, First-principles studies of the Gilbert damping and exchange interactions for half-metallic Heuslers alloys, Phys.
-Rev.
-B 93, 214439 (2016).
-- Kuroda et al.
-[2020] F.
-Kuroda, T.
-Fukushima, and T.
-Oguchi, First-principles study of magnetism and phase stabilities of V2 based antiferromagnetic Heusler alloys, Journal of Applied Physics 127, 193904 (2020).
-- Skaftouros et al. [2013] S. Skaftouros, K. Özdoğan, E. Şaşıoğlu, and I. Galanakis, Generalized Slater-Pauling rule for the inverse Heusler compounds, Phys. Rev. B 87, 024420 (2013).
-- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
-- Kresse and Joubert [1999] G. Kresse and D. Joubert, From ultrasoft pseudopotentials to the projector augmented-wave method, Phys. Rev. B 59, 1758 (1999).
-- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
-- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
-- Hohenberg and Kohn [1964] P. Hohenberg and W. Kohn, Inhomogeneous electron gas, Phys. Rev. 136, B864 (1964).
-- Kohn and Sham [1965] W. Kohn and L. J. Sham, Self-consistent equations including exchange and correlation effects, Phys. Rev. 140, A1133 (1965).
-- Li et al.
-[2016] P.
-Li, X.
-Liu, M.
-Chen, P.
-Lin, X.
-Ren, L.
-Lin, C.
-Yang, and L.
-He, Large-scale ab initio simulations based on systematically improvable atomic basis, Computational Materials Science 112, 503 (2016), computational Materials Science in China.
-- Zhou et al.
-[2025] W.
-Zhou, D.
-Zheng, Q.
-Liu, D.
-Lu, Y.
-Liu, P.
-Lin, Y.
-Huang, X.
-Peng, J.
-J.
-Bao, C.
-Cai, Z.
-Jin, J.
-Wu, H.
-Zhang, G.
-Jin, Y.
-Ji, Z.
-Shen, X.
-Liu, L.
-Sun, Y.
-Cao, M.
-Sun, J.
-Liu, T.
-Chen, R.
-Liu, Y.
-Li, H.
-Han, X.
-Liang, T.
-Bao, N.
-Chen, H.
-Ren, X.
-Zhang, Z.
-Liu, Y.
-Fu, M.
-Liu, Z.
-Li, T.
-Wen, Z.
-Tang, Y.
-Xu, W.
-Duan, X.
-Wang, Q.
-Gu, F.-Z.
-Dai, Q.
-Zheng, J.
-Zhao, Y.
-Zhang, Q.
-Ou, H.
-Jiang, S.
-Liu, B.
-Xu, S.
-Xu, X.
-Ren, L.
-He, L.
-Zhang, and M.
-Chen, ABACUS: An electronic structure analysis package for the AI era, arXiv e-prints , 2501.08697 (2025).
-- Schlipf and Gygi [2015] M. Schlipf and F. Gygi, Optimization algorithm for the generation of ONCV pseudopotentials, Computer Physics Communications 196, 36 (2015).
-- He et al. [2021] X. He, N. Helbig, M. J. Verstraete, and E. Bousquet, TB2J: A python package for computing magnetic interaction parameters, Computer Physics Communications 264, 107938 (2021).
-- Skubic et al.
-[2008] B.
-Skubic, J.
-Hellsvik, L.
-Nordström, and O.
-Eriksson, A method for atomistic spin dynamics simulations: implementation and examples, Journal of Physics: Condensed Matter 20, 315203 (2008).
-- Goraus et al.
-[2023] J.
-Goraus, M.
-Fijałkowski, J.
-Czerniewski, and W.
-Gumulak, Magnetic properties of V2MnGa, V2MnAl, V2FeGa and V2FeAl, Journal of Magnetism and Magnetic Materials 585, 171103 (2023).
-- Zhang et al.
-[2012] X.
-Zhang, X.
-Dai, G.
-Chen, H.
-Liu, H.
-Luo, Y.
-Li, W.
-Wang, G.
-Wu, and G.
-Liu, First-principle study for full-Heusler compounds V2YAl (Y=V, Cr, Mn, Fe, Co, Ni) and a discussion to Slater–Pauling rule, Computational Materials Science 59, 1 (2012).
-- Faleev et al.
-[2017] S.
-V.
-Faleev, Y.
-Ferrante, J.
-Jeong, M.
-G.
-Samant, B.
-Jones, and S.
-S.
-P.
-Parkin, Unified explanation of chemical ordering, the Slater-Pauling rule, and half-metallicity in full Heusler compounds, Phys.
-Rev.
-B 95, 045140 (2017).
-- Xu et al.
-[2024] Y.
-Xu, Y.-C.
-Wang, X.
-Jin, H.
-Liu, Y.
-Liu, H.
-Song, and F.
-Tian, Mechanism of magnetic phase transition in correlated magnetic metal: insight into itinerant ferromagnet Fe3‑δGeTe2, Commun.
-Phys.
-7, 381 (2024).
-- Xu et al. [2025a] Y. Xu, X. Jin, J. Xiang, H. Zhang, and F. Tian, investigation of the magnetic phase transition in the itinerant ferromagnet , Phys. Rev. B 111, 155142 (2025a).
-- Xu et al. [2025b] Y. Xu, X. Jin, H. Tang, and F. Tian, Electronic structures and magnetism in van der waals flat-band material , Europhysics Letters 151, 36002 (2025b).
+Sci.
+59, pp.
+1–5.
+External Links: ISSN 0927-0256, Document, Link Cited by: §II, §III.1, §III.2, Table 2, Table 2.
+- [56] (2025-01) ABACUS: an electronic structure analysis package for the AI era. arXiv, pp. 2501.08697. External Links: Document Cited by: §II.
+- [57] (2022) Structural and magnetic properties of inverse-Heusler Mn2FeSi alloy powder prepared by ball milling.
+Materials 15 (3), pp.
+697.
+External Links: Link, ISSN 1996-1944, Document Cited by: §I.
+- [58] (2004-04) Spintronics: fundamentals and applications. Rev. Mod. Phys. 76, pp. 323–410. External Links: Document, Link Cited by: §I.